A proton moving at 6.60 106 m/s through a magnetic field of magnitude 1.78 T experiences a magnetic force of magnitude 7.20 10-13 N. What is the angle between the proton's velocity and the field? (Enter both possible answers from smallest to largest. Enter only positive values between 0 and 360.)
Answer #1


v = 6.6*10^6 m/s, F =7.2*10^-13 N, B = 1.78 T

proton charge q = 1.6*10^-19 C

Magnetic force F = q(vxB)

F = qvB sin(theta)

7.2*10^-13 = 1.6*10^-19*6.6*10^6*1.78*sin(theta)

theta =22.52 degrees , 157.5 degrees